# Shifting the Lens: Detecting Malicious npm Packages using Large Language Models

### v2

@@ VIII Related Work
Malicious Package Detection: Several approaches have been proposed to detect malicious packages that can be classified as rule-based [14, 57, 17, 19], differential analysis [35, 45, 47, 41] and machine learning [34, 42] approaches. Rule-based approaches often rely on predefined rules about package metadata (e.g., package name), suspicious imports, and method calls. Many studies [14, 57, 17, 19] have focused on detecting static features like libraries that access the network, file system access, features in package metadata, dynamic analysis, or language-independent features like commit logs and repository metadata to flag packages as suspicious using heuristic rules. Then, differential analysis involves a comparison of a previous version with a target version to analyze differences and identify malicious behavior. These tools [35, 45, 47, 41] leverage artifacts, file hashes, and“phantom lines” to show differences between versions of packages and source code repositories. The studies [34, 42] evaluated various supervised ML models to identify the effective ML models to detect malware. Additionally, multiple research studies [47, 44] exist on detecting typosquat attacks. Despite such vast research in this domain, Ohm et al.  [37] revealed in a Systemization of Knowledge (SOK) study that none of the presented approaches can be considered a complete and isolated solution to detect malicious behavior. Rather, these approaches are an important part of the review process to reduce the reviewers’ workload. Ohm et al. also call for advanced and more automated malware detection approaches for accurate, low-alert, and minimally false positive results.
@@
Large Language Model (LLMs): LLMs are now widely applied in software engineering across various tasks, including code generation, testing, program repair, and summarization [6, 55]. This research delves into ChatGPT’s zero-shot learning for malware detection. Zero-shot is a technique that enables models to tackle unseen tasks without prior examples. Zero-shot learning has shown promising results in various NLP tasks [10, 50], with Chain-of-thought (CoT) or step-by-step reasoning prompts enhancing LLMs’ ability to produce intermediate reasoning steps [50]. Studies [54, 25, 26] have identified effective CoT components, such as zero-shot CoT and role-play prompting, that improve reasoning quality. Kojima et al.[25] demonstrated LLMs’ proficiency in zero-shot reasoning, while Kong et al. [26] found that role-play prompting yields more effective CoT. Madaan et al. [33] observed that the SELF-REFINE method, applied to tasks ranging from dialog to reasoning using GPT-3.5 and GPT-4, significantly improves performance, validated by human and automatic metrics. Zelikman et al. [58] confirmed the utility of self-generated rationales for model self-improvement.
@@
White et al. [52, 51] demonstrated that augmenting the model with domain-specific information improves the model’s performance since language models find more value in detecting specific identifiers in code structures when generating code summaries. Chen et al. [11] investigated the performance of GPT-3 code generation. They explored the model’s ability to understand and generate code, offering valuable insights into the potential of using such models for code-related tasks. Ahmed et al. [6] showed that the GPT model can significantly surpass state-of-the-art models for code summarization, leveraging project-specific training. In another study, Sobania et al. [43] utilized ChatGPT for code bug fixing, further improving the success rate of bug fixing. Purba et al. [39] conducted a study measuring LLMs performance in detecting software vulnerabilities. The authors concluded that LLMs do not perform well in detecting software vulnerabilities, but models are particularly good at finding common patterns that are involved in vulnerable code. Hamer et al. [21] compare the security vulnerabilities of code from ChatGPT and StackOverflow, finding that ChatGPT-generated code has 20% fewer vulnerabilities and fewer CWE types than SO.
IX Threats To Validity
We used the MalwareBench dataset [56] (Section II-A), which contains past malicious samples from prior research [36, 14, 20, 27]. However, the dataset may contain bias toward a certain type of malicious code, and the repetition of the same attack across multiple packages might have skewed the dataset. As discussed in Section VII, model performance might differ with complex attacks or large files where malware spreads in different forms all over the file. Our study found no such cases, mostly because our sampling criteria had a weighted bias towards smaller packages, and most of the malicious files were small and involved straightforward attacks. Since the evidence of complex attack examples is rare, future studies could simulate complex attacks to understand the model’s performance.
While the results obtained in evaluating our workflow are promising, the false negative result and non-deterministic nature of LLMs raise concerns about model performance, especially misclassification due to the lack of diversity and ambiguous and noisy output. Additionally, due to ChatGPT’s continual evolution, our specific results may not be replicable in future iterations. However, the proposed workflow can be applied to future studies, ensuring their relevance and applicability despite technological changes. Additionally, the authors [56] of the MalwareBench dataset only vetted 900 packages out of 10k neutral packages and assumed others as neutral. Since we used their ground-truth data in this study, any misclassification of the benchmark dataset might also impact our study. We did our best to check that our dataset’s neutral samples do not include malicious code by manually evaluating 2,530 files out of 20,926 files. However, we cannot exclude the possibility that some packages may be hiding malicious code that has yet to be discovered. Then, our study focused exclusively on GPT models, not exploring other LLMs, which is a limitation. Adapting our workflow for other models would necessitate a similar, if not greater, investment of time tailored to the model’s unique attributes. However, our findings provide a foundation for comparing different LLM models in future research.
X Conclusion and Future Study
@@
In this work, we present SecurityAI, a malicious code review workflow, and investigate the feasibility of LLMs in code review to detect malicious npm packages. Our study can be used as a baseline study in the security domain. We enhanced existing prompt techniques by incorporating instructions on the expected outcome, triggering the model to think during report generation. We conducted a baseline comparison of the GPT-3 and GPT-4 models with a commercial static analysis tool, which showed promising results for the GPT models with low misclassification alert rates. SecurityAI detected 2,057 malicious packages (both models) out of 2,180 packages, presenting an approach that reduces the burden of manual review on security analysts and offers findings with minimal misclassification (332). Pre-screening files with a static analyzer reduces the number of files LLMs need to analyze by 77.9% and significantly cuts costs for GPT-3 and GPT-4 by 60.9% and 76.1%, respectively. Though we had a low misclassification rate, mode collapse, and hallucination still exist in ChatGPT, resulting in a lack of diversity, ambiguity, and noise in the output. Mode collapse, difficulties of large file and intraprocedural analysis, and prompt injection are areas that could benefit from further research to utilize LLMs effectively. Researchers can use SecurityAI workflow and prompt used in this study for future research to evaluate other LLMs and assess their performance across more diverse datasets.

### v3

@@ VIII Related Work
Malicious Package Detection: Several approaches have been proposed to detect malicious packages that can be classified as rule-based [9, 56, 12, 14], differential analysis [31, 43, 45, 38] and machine learning [30, 39] approaches. Rule-based approaches often rely on predefined rules about package metadata (e.g., package name), suspicious imports, and method calls. Many studies [9, 56, 12, 14] have focused on detecting static features like libraries that access the network, file system access, features in package metadata, dynamic analysis, or language-independent features like commit logs and repository metadata to flag packages as suspicious using heuristic rules. Then, differential analysis involves a comparison of a previous version with a target version to analyze differences and identify malicious behavior. These tools [31, 43, 45, 38] leverage artifacts, file hashes, and“phantom lines” to show differences between versions of packages and source code repositories. The studies [30, 39] evaluated various supervised ML models to identify the effective ML models to detect malware. Yu et al. [54] used LLM to translate malicious functions from other languages (e.g., C/C++, Python, and Go) into JavaScript. Singla et al. [40] used GPT and Bard models to characterize software supply chain failure reports/blogs into the type of compromise, intent, nature, and impact. Additionally, multiple research studies [45, 42] exist on detecting typosquat attacks. Despite such vast research in this domain, Ohm et al.  [33] revealed in a Systemization of Knowledge (SOK) study that none of the presented approaches can be considered a complete and isolated solution to detect malicious behavior. Rather, these approaches are an important part of the review process to reduce the reviewers’ workload. Ohm et al. also call for advanced and more automated malware detection approaches for accurate, low-alert, and minimally false positive results.
@@
Large Language Model (LLMs): LLMs are now widely applied in software engineering across various tasks, including code generation, testing, program repair, and summarization [5, 53]. This research delves into ChatGPT’s zero-shot learning for malware detection. Zero-shot is a technique that enables models to tackle unseen tasks without prior examples. Zero-shot learning has shown promising results in various NLP tasks [7, 48], with Chain-of-thought (CoT) or step-by-step reasoning prompts enhancing LLMs’ ability to produce intermediate reasoning steps [48]. Studies [52, 20, 21] have identified effective CoT components, such as zero-shot CoT and role-play prompting, that improve reasoning quality. Kojima et al.[20] demonstrated LLMs’ proficiency in zero-shot reasoning, while Kong et al. [21] found that role-play prompting yields more effective CoT. Madaan et al. [28] observed that the SELF-REFINE method, applied to tasks ranging from dialog to reasoning using GPT-3.5 and GPT-4, significantly improves performance, validated by human and automatic metrics.
@@
White et al. [50, 49] demonstrated that augmenting the model with domain-specific information improves the model’s performance since language models find more value in detecting specific identifiers in code structures when generating code summaries. Ahmed et al. [5] showed that the GPT model can significantly surpass state-of-the-art models for code summarization, leveraging project-specific training. In another study, Sobania et al. [41] utilized ChatGPT for code bug fixing, further improving the success rate of bug fixing. Purba et al. [36] conducted a study measuring LLMs performance in detecting software vulnerabilities. Hamer et al. [16] compare the security vulnerabilities of code from ChatGPT and StackOverflow, finding that ChatGPT-generated code has 20% fewer vulnerabilities and fewer CWE types than SO.
IX Conclusion and Future Study
@@
In this work, we present SocketAI, a malicious code review workflow, and investigate the feasibility of LLMs in code review to detect malicious npm packages. Our study can be used as a baseline study in the security domain. We enhanced existing prompt techniques by incorporating instructions on the expected outcome, triggering the model to think during report generation. We conducted a baseline comparison of the GPT-3 and GPT-4 models with a commercial static analysis tool, which showed promising results for the GPT models with low misclassification alert rates. SocketAI detected 2,057 malicious packages (both models) out of 2,180 packages, presenting an approach that reduces the burden of manual review on security analysts and offers findings with minimal misclassification (332). Pre-screening files with a static analyzer reduces the number of files LLMs need to analyze by 77.9% and significantly cuts costs for GPT-3 and GPT-4 by 60.9% and 76.1%, respectively. Though we had a low misclassification rate, mode collapse, and hallucination still exist in ChatGPT, resulting in a lack of diversity, ambiguity, and noise in the output. Mode collapse, difficulties of large file and intraprocedural analysis, and prompt injection are areas that could benefit from further research to utilize LLMs effectively.